### v1

@@ Appendix G Multiple scattering and time-step
The probability not to scatter if we have only one time-step over an interval $\Delta t$ is
@@
Given that we have $n$ equally spaced time-steps of size $\Delta t/n$, it becomes (because of Bernoulli’s inequality)
@@
Here $P_{\mathrm{nsm}}>P_{\mathrm{ns}}$ for $n>1$. Hence, smaller time-steps slightly increase the number of unscattered particles. The probability to scatter once can be expressed as
@@
Given that we have $n$ equally spaced time-steps it becomes
@@
Here $P_{\mathrm{osm}}<P_{\mathrm{os}}$ for $n>1$. Hence, smaller time-steps slightly decrease the number of particles that have scattered only once. Furthermore, there is a recent study by Fischer et al. (2024a) that gives the limit for $n\to\infty$. We note that these equations are not straightforwardly applicable to the simulations, for example, a numerical particle can interact with multiple neighbours per time-step and $N_{\mathrm{ngb}}$ is not taken into account in the equations. However, the equations provide a qualitative understanding and thus provide insight into the limitations of the simulations.
@@
To illustrate these limitations, we show in Fig. 17 the fixed-angle deflection test simulated with a varying number of time-steps for a given time $t$. More precisely, we simulate this deflection test with a fixed angle of $\theta_{0}=0.2$ for a reference time-step of $\Delta t=0.0012$ Gyr and other equally spaced time-steps of $\Delta t\in\{0.1536,0.0768,0.0384,0.0192,0.0096,0.0048,0.0024\}$ Gyr. We use the same simulation set-up described in Sect. 5.1, a transfer cross section of $\sigma_{\mathrm{T}}/m_{\chi}=100\,{\rm cm}^{2}\,{\rm g}^{-1}$ and $N_{\mathrm{ngb}}=16$ neighbouring particles. The angle distribution is the (already normalised) mean value of $8$ simulations employing different seeds for the pseudo-random number generator.
@@
The three panels of Fig. 17 show the distribution of the deflection angle at different times. The first peak in the plot (from left to right) corresponds to the unscattered particles, the second peak to the particles that have scattered once, and the third peak is mainly particles that have scattered twice. The height of the first peak decreases over the course of the simulation, while the height of the second peak first increase and later decreases. It is clearly visible that the height of the peaks depends on the time resolution, that is how many time-steps we use to arrive at the solution. In line with the previous estimates, we find that the number of particles that have not scattered yet increases with increasing time resolution. In contrast, for the particles that have scattered once the picture is more complicated, in an initial phase a better time resolution leads to a smaller peak while at a later stage, it leads to a higher peak. At that late time also the peak height is decreasing with time. We note that these findings are in line with Eq. (75).
@@
For each of our simulation times (set by multiples of the largest time-step), $t\in\{0.1536,0.3072,0.4608,0.6144\}$ Gyr we can compute the ratio of peak heights between the different time resolutions and the reference ($\Delta t=0.0012$ Gyr) for both peaks to better see the multiple scattering effect. We show the result for the first (second) peak in the left (right) panel of Fig. 18 as a function of time. Again, we see from the first peak that the number of time-steps we use clearly influences the number of unscattered particles. For both peaks, we find that the relative difference is increasing with time at the later times we show here.
@@
We can conclude that one needs time-steps not only small enough to ensure $P_{\mathrm{os}}<1$ but $P_{\mathrm{os}}\ll 1$ to obtain accurate results, because numerically we can only describe the single scatterings but not the multiple ones of physical particles. How important this is for a simulation, depends on how crucial resolving the precise angular dependence for the evolution of the SIDM distribution is. While we have chosen a sensitive test, this problem might be less relevant for typical astrophysical simulations. In consequence, the optimal choice of the time-step, to achieve a good trade-off between accuracy and performance might be problem-dependent. Thus the value for $\kappa$ in the time-step criterion (Eq. (43)), which effectively limits the interaction probability, could be chosen depending on the problem at hand.
@@
Figure 17: Fixed-angle deflection test for several time resolutions. The distribution of the deflection angles for runs using a different number of time-steps to reach the solution is shown. The three panels give the results for different points in time.
@@
Figure 18: Fixed-angle deflection test with a varying number of time-steps to the solution. The ratio of the number of particles in the first peak (i.e. particles that have not yet scattered) for the various time resolutions compared to the reference simulation is displayed as a function of time in the left panel. The results for the second peak (i.e. particles that have scattered once) are shown in the right panel.

### v2

@@ Appendix G Multiple scattering and time step
The probability not to scatter if we have only one time step over an interval $\Delta t$ is
@@
Given that we have $n$ equally spaced time steps of size $\Delta t/n$, it becomes (because of Bernoulli’s inequality)
@@
Here $P_{\mathrm{nsm}}>P_{\mathrm{ns}}$ for $n>1$. Hence, smaller time steps slightly increase the number of unscattered particles. The probability to scatter once can be expressed as
@@
Given that we have $n$ equally spaced time steps, it becomes
@@
Here $P_{\mathrm{osm}}<P_{\mathrm{os}}$ for $n>1$. Hence, smaller time steps slightly decrease the number of particles that have scattered only once. Furthermore, there is a recent study by Fischer et al. (2024a) that gives the limit for $n\to\infty$. We note that these equations are not straightforwardly applicable to the simulations, for example, a numerical particle can interact with multiple neighbours per time step and $N_{\mathrm{ngb}}$ is not taken into account in the equations. However, the equations provide a qualitative understanding and thus provide insight into the limitations of the simulations.
@@
To illustrate these limitations, we show in Fig. 17 the fixed-angle deflection test simulated with a varying number of time steps for a given time $t$. More precisely, we simulate this deflection test with a fixed angle of $\theta_{0}=0.2$ for a reference time step of $\Delta t=0.0012$ Gyr and other equally spaced time steps of $\Delta t\in\{0.1536,0.0768,0.0384,0.0192,0.0096,0.0048,0.0024\}$ Gyr. We use the same simulation set-up described in Sect. 5.1, a transfer cross section of $\sigma_{\mathrm{T}}/m_{\chi}=100\,{\rm cm}^{2}\,{\rm g}^{-1}$ and $N_{\mathrm{ngb}}=16$ neighbouring particles. The angle distribution is the (already normalised) mean value of $8$ simulations employing different seeds for the pseudo-random number generator.
@@
The three panels of Fig. 17 show the distribution of the deflection angle at different times. The first peak in the plot (from left to right) corresponds to the unscattered particles, the second peak to the particles that have scattered once, and the third peak is mainly particles that have scattered twice. The height of the first peak decreases over the course of the simulation, while the height of the second peak first increase and later decreases. It is clearly visible that the height of the peaks depends on the time resolution, that is, how many time steps we use to arrive at the solution. In line with the previous estimates, we find that the number of particles that have not scattered yet increases with increasing time resolution. In contrast, for the particles that have scattered once the picture is more complicated, in an initial phase a better time resolution leads to a smaller peak while at a later stage, it leads to a higher peak. At that late time also the peak height is decreasing with time. We note that these findings are in line with Eq. (75).
@@
For each of our simulation times (set by multiples of the largest time step), $t\in\{0.1536,0.3072,0.4608,0.6144\}$ Gyr we can compute the ratio of peak heights between the different time resolutions and the reference ($\Delta t=0.0012$ Gyr) for both peaks to better see the multiple scattering effect. We show the result for the first (second) peak in the left (right) panel of Fig. 18 as a function of time. Again, we see from the first peak that the number of time steps we use clearly influences the number of unscattered particles. For both peaks, we find that the relative difference is increasing with time at the later times we show here.
@@
We can conclude that one needs time steps not only small enough to ensure $P_{\mathrm{os}}<1$ but $P_{\mathrm{os}}\ll 1$ to obtain accurate results, because numerically we can only describe the single scatterings but not the multiple ones of physical particles. How important this is for a simulation, depends on how crucial resolving the precise angular dependence for the evolution of the SIDM distribution is. While we have chosen a sensitive test, this problem might be less relevant for typical astrophysical simulations. In consequence, the optimal choice of the time step to achieve a good trade-off between accuracy and performance might be problem dependent. Thus the value for $\kappa$ in the time step criterion (Eq. (43)), which effectively limits the interaction probability, could be chosen depending on the problem at hand.
@@
Figure 17: Fixed-angle deflection test for several time resolutions. The distribution of the deflection angles for runs using a different number of time steps to reach the solution is shown. The three panels give the results for different points in time.
@@
Figure 18: Fixed-angle deflection test with a varying number of time steps to the solution. The ratio of the number of particles in the first peak (i.e. particles that have not yet scattered) for the various time resolutions compared to the reference simulation is displayed as a function of time in the left panel. The results for the second peak (i.e. particles that have scattered once) are shown in the right panel.